# Black Hole Waterfall: a unitary phenomenological model for black hole evaporation with Page curve

### v1

@@ Appendix A Derivation of One Shot Black Hole Waterfall model
where we have “summed down” each column labeled by $d$, and introduced the collective total-emission variables $k_{d}\equiv\sum_{i=1}^{N}n_{id}$, representing the total number of particles emitted by idler pump $i_{d-1}$ in $N$ times steps (where $i_{0}\equiv p$ is the BH pump). Approximating all the denominators in each $P_{n_{emit}}^{(n_{initial})}$ in Eq.(67) by unity gives the collective unnormalized states signal/idler states $|\tilde{\Phi}^{(N)}_{k_{1},k_{2},\ldots,k_{D}}\rangle$. These states contain the rest of the summations over the remaining collective variables defined as $j_{i,d}\leavevmode\nobreak\ \equiv\leavevmode\nobreak\ \sum_{i^{\prime}=0}^{i}n_{i^{\prime},d}$, generalizing the sums int the state $|\tilde{\Phi}^{(N)}_{k_{1}}\rangle$ in Eq.(22) for the depth $D=1$ case. As in the original $D=1$ One Shot method, we can create a finer approximation to these denominators by using their lowest power of $z$ (since we assume $z\ll 1$), and approximate them by their largest contributions. This yields the factor of $\frac{1}{(1-z^{k_{d-1}-k_{d}+1})^{N-d}}$ in Eq.(55). Essentially, the factor $\frac{1}{(1-z^{{n_{p0}}-k_{1}+1})^{N-d}}$ at depth $d=1=D$ goes to $\frac{1}{(1-z^{k_{d-1}-k_{d}+1})^{N-d}}$ at intermediate depth $d\in[1,D]$, with one caveat. The former factor can be pulled all the way to the left (in front of $\sum_{k_{1}}^{{n_{p0}}}$) in the nested summations $\sum_{k_{1}}^{{n_{p0}}}\sum_{k_{1}}^{k_{2}}\sum_{k_{2}}^{k_{3}}\cdots\sum_{k_{D}}^{k_{D-1}}$ over the collective emission variables $(k_{1},k_{2},k_{3},\ldots,k_{D})$, since it only depends on $k_{1}$. However, subsequent factors $\frac{1}{(1-z^{k_{d-1}-k_{d}+1})^{N-d}}$ for the idler pump at intermediate depth $d$ can only be pulled as far to left as the summation $\sum_{k_{d}}^{k_{d-1}}$ (hence the power $N-d$ now instead of $N-1$) [46].

### v2

@@ Appendix A Derivation of One Shot Black Hole Waterfall model
where we have “summed down” each column labeled by $d$, and introduced the collective total-emission variables $k_{d}\equiv\sum_{i=1}^{N}n_{id}$, representing the total number of particles emitted by idler pump $i_{d-1}$ in $N$ times steps (where $i_{0}\equiv p$ is the BH pump). Approximating all the denominators in each $P_{n_{emit}}^{(n_{initial})}$ in Eq.(67) by unity gives the collective unnormalized states signal/idler states $|\tilde{\Phi}^{(N)}_{k_{1},k_{2},\ldots,k_{D}}\rangle$. These states contain the rest of the summations over the remaining collective variables defined as $j_{i,d}\leavevmode\nobreak\ \equiv\leavevmode\nobreak\ \sum_{i^{\prime}=0}^{i}n_{i^{\prime},d}$, generalizing the sums int the state $|\tilde{\Phi}^{(N)}_{k_{1}}\rangle$ in Eq.(22) for the depth $D=1$ case. As in the original $D=1$ One Shot method, we can create a finer approximation to these denominators by using their lowest power of $z$ (since we assume $z\ll 1$), and approximate them by their largest contributions. This yields the factor of $\frac{1}{(1-z^{k_{d-1}-k_{d}+1})^{N-d}}$ in Eq.(55). Essentially, the factor $\frac{1}{(1-z^{{n_{p0}}-k_{1}+1})^{N-d}}$ at depth $d=1=D$ goes to $\frac{1}{(1-z^{k_{d-1}-k_{d}+1})^{N-d}}$ at intermediate depth $d\in[1,D]$, with one caveat. The former factor can be pulled all the way to the left (in front of $\sum_{k_{1}}^{{n_{p0}}}$) in the nested summations $\sum_{k_{1}}^{{n_{p0}}}\sum_{k_{1}}^{k_{2}}\sum_{k_{2}}^{k_{3}}\cdots\sum_{k_{D}}^{k_{D-1}}$ over the collective emission variables $(k_{1},k_{2},k_{3},\ldots,k_{D})$, since it only depends on $k_{1}$. However, subsequent factors $\frac{1}{(1-z^{k_{d-1}-k_{d}+1})^{N-d}}$ for the idler pump at intermediate depth $d$ can only be pulled as far to left as the summation $\sum_{k_{d}}^{k_{d-1}}$ (hence the power $N-d$ now instead of $N-1$)  [^6] Formally we should use the power $N-d$ if it is non-negative, and use, say $N-1$ if $N-d<0$, since the (discrete) time steps are $N\geq 1$. However, since $z<<1$ we have found that numerically there is not much effect in the early time steps if we just continue to use the power $N-d$. A conditional “if-else” coding statement, such as $\textrm{If}[N-d\geq 0,N-d,N-1]$, tends to slow down the numerical computations..